# PixFoundation: Are We Heading in the Right Direction with Pixel-level Vision Foundation Models?

### v2

@@ Abstract
Multiple works have emerged to push the boundaries on multi-modal large language models (MLLMs) towards pixel-level understanding. Such approaches have shown strong performance on benchmarks for referring expression segmentation and grounded conversation generation. The current trend in pixel-level MLLMs is to train with pixel-level grounding supervision on large-scale labelled data. However, we show that such MLLMs when evaluated on recent challenging vision centric benchmarks, exhibit a weak ability in visual question answering. Surprisingly, some of these methods even downgrade the grounding ability of MLLMs that were never trained with such supervision. In this work, we propose two novel challenging benchmarks and show that MLLMs without pixel-level grounding supervision can outperform the state of the art in such tasks when evaluating both the pixel-level grounding and visual question answering. We propose simple baselines to extract the grounding information that can be plugged into any MLLM, which we call as PixFoundation. More importantly, we study the research question of “When does grounding emerge in MLLMs that are not trained with pixel-level grounding supervision?” We show that grounding can coincide with object parts or location/appearance information. Code repository is at https://github.com/MSiam/PixFoundation/.
@@ 1 Introduction
There have been numerous advancements in pixel-level image and video understanding, including tasks such as image/video segmentation [29, 14, 8, 17], pixel-level visual grounding and reasoning [16, 9], depth estimation [23] and tracking [21]. The majority of these have been transformed with the emergence of foundation models [1], specifically multi-modal large language models (MLLMs) [11, 3]. Nonetheless, pixel-level MLLMs have shown degradation in their capabilities in chat performance [9]. Recent models tried to address this gap [27, 26], yet they relied on standard evaluation benchmarks overlooking the shortcomings of current MLLMs.
@@
Recent efforts explored the shortcomings of MLLMs in vision centric benchmarks [19, 18]. Such benchmarks focused on challenging visual patterns and tasks such as counting and relative positioning. Nonetheless, these benchmarks did not evaluate the recent pixel-level MLLMs. In this work we propose challenging vision centric benchmarks that are dedicated to evaluate these models. Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with full grounding supervision heading in the right direction to improve both grounding and visual question answering (VQA)?”. Our findings show that the majority of pixel-level MLLMs still fall short in such a challenging setting. While evidently, some of these show superior performance in visual grounding, we show that MLLMs that were not trained with pixel-level grounding can have better performance without degrading their chat capabilities.
There have been recent works showing training-free open vocabulary segmentation emerging from vision language models [20, 13, 5]. Concurrent work has specifically explored emerging grounding in MLLMs [2]. We are inspired by the previous work to establish a baseline on our proposed benchmarks. However, unlike their method we focus on the second research question of “When does grounding emerge in MLLMs that are not trained with full grounding supervision?”. Previous efforts studying emergent segmentation and/or grounding assumed correspondence with the specific language tokens (i.e., nouns or noun phrases) of the objects of interest. However, our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the object. We show the most frequent emergence in PixMMVP happens in the last 40-60% of the output, and can coincide with concepts about object parts, position, color or context of these objects. Fig. 1 summarizes our research questions.
@@
In summary, our contributions include: (i) Proposing pixel-level vision centric benchmarks, PixMMVP and PixCV-Bench, with segmentation annotations and referring expression of the object of interest in the corresponding questions [18, 19]. (ii) Benchmarking recent efforts in pixel-level MLLMs where we show that they degrade VQA capabilities. More importantly, some of them lag in visual grounding with respect to simple techniques of extracting the segmentation from vanilla MLLMs, i.e., MLLMs that are not trained for pixel-level grounding. (iii) We provide a simple mechanism for extracting segmentation from vanilla MLLMs, with an understanding of when grounding emerges. Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest. We call it PixFoundation, since it can be plugged with any MLLM to mine for pixel-level understanding.
@@ 2 Related work
Pixel-level Vision Foundation Models. There have been various vision foundation models released that were trained with supervision for the segmentation task (e.g., SAM, SAM 2.0) [8, 17]. Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision [13, 5, 20]. Yet, they relied on earlier foundation models that did not incorporate the power of large language models. Combining large language models with vision has been extensively researched with pioneering works such as Llava [11, 12] and instruct-BLIP [3]. Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with full supervision [9, 16, 26, 27, 26, 27]. However, these methods were lagging in their chat performance. Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs [19, 18]. Hence, it is still unclear if the pixel-level grounding supervision helped to improve their ability on these challenging tasks or not. In this work, we focus on the previous question to have a better understanding of their performance. Concurrent work, has shown that without pixel-level supervision there is an emerging ability to perform pixel-level grounding [2]. We rely on this method as our baseline, but unlike previous works we provide an insight on when grounding emerges in such MLLMs. We propose another baseline that uses a novel and simple mechanism to perform mask selection while taking the previous insight into consideration. Our baseline outperforms the previous methods trained with full grounding supervision and the ones without it.
Benchmarking Multi-modal Large Language Models. There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU [25]) and pixel-level benchmarks (e.g., refCOCO/+/g [24, 7]). These have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding. Nonetheless, there have been various works that discussed the shortcomings of MLLMs. One of them discussed the shortcomings in CLIP [15], which is used in various MLLMs as a visual backbone. They proposed a benchmark, MMVP [19], that is focused on the visual aspects within a VQA task. More recently, CV-Bench [18] focused on two major tasks that are vision focused which are counting and relative positioning. Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output. Nonetheless, they still provide quite challenging scenarios that can act as a strong benchmark for the pixel-level MLLMs counterpart. In this work, we extend these two benchmarks to augment it with pixel-level annotations and referring expressions that correspond to the object of interest within their VQA task. These help to understand if failures occurring in these benchmarks stem from grounding or VQA capabilities.
@@ 3 Method and benchmarks
In this section, we describe our two benchmarks and probing techniques for pixel-level MLLMs and MLLMs that were not trained with pixel-level grounding supervision.
3.1 Benchmarks
@@
PixMMVP benchmark: We build upon the recently released MMVP [19] which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images. We augment the aforementioned dataset and manually annotate each question with the corresponding object of interest referring expression, e.g. an elderly person or the butterfly’s feet. There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded. Examples include questions inquiring on the view direction of the camera which are not tied to a specific entity. Our manual referring expression annotations are as fine-grained as possible. These expressions correspond to what needs to be grounded in the image to answer the question. Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator [4].
@@
PixCV-Bench benchmark: For this benchmark we build upon the 2D component of the recently released CV-Bench [18]. We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K [28] and COCO [10]), which can be used in our proposed benchmark. However, the publicly released CV-Bench does not identify the objects in question and their corresponding segmentation. As such we use GPT-4o to parse the questions and identify the objects of interest automatically, followed by manual inspection and correction. Specifically, we collect the classes in each image from the corresponding dataset and construct a list of class choices “1. CLS1, 2. CLS2, …”. Then we prompt GPT-4o with the following, “Provide number only as an answer. Identify the objects of interest in the following question: QUESTION ? 1. CLS1, 2. CLS2, … ”. This provides us with the categories per question that highlights the objects of interest. While seemingly these are categorical annotations not referring expressions, certain scenarios in CV-Bench are different. Specifically, in the relative positioning task all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categorical annotations.
@@
Afterwards, we use the selected categories from GPT-4o to retrieve the corresponding segmentation mask/s per image. Furthermore, we use a custom annotation tool to manually filter the objects in the question, e.g. selecting only the object mask annotated by the red box when referred to it and filtering out the other instances for that same class. Another example that needs manual filtration when the class in question is a broader category than what is inquired upon, e.g., “Pendant Lamp” which is under the category of “Lamp” in ADE20K. In such a case, we filter out the masks of other types such as “Table Lamp”. Moreover, we identify missing annotations in rare occasions that require additional intervention and manually annotate these missing objects. We provide the final PixCV-Bench with referring expressions and their segmentation annotations that can be used to evaluate the grounding ability in relation to the original VQA task. Appendix A provides visual examples from our benchmarks.
3.2 A Pixel-level MLLMs study
We utilize the two proposed benchmarks, PixMMVP and PixCV-Bench, to evaluate how the current trend in pixel-level MLLMs that relies on training with grounding supervision perform in such challenging tasks. Furthermore, we inspect the failures of these pixel-level MLLMs and explore simple approaches to pixel-level understanding from MLLMs that overcome the previous shortcomings.
@@ 3.2 A Pixel-level MLLMs study
Pixel-level MLLMs shortcomings. We highlight the failures for the current state-of-the-art pixel-level MLLMs through three probing techniques. First, we highlight the degraded performance in VQA from most of the pixel-level MLLMs that are trained with pixel-level grounding supervision. We use for that the following prompt, “IMGQUESTION? OPTION1 OPTION2…”, as shown in Figure 2. Certain pixel-level MLLMs tend to answer the aforementioned question while outputting a corresponding segmentation mask/s for the objects of interest. Notably, the worst two models in this task, LISA [9] and GLAMM [16], are not able to provide an answer and rather refer to a segmentation mask. On the other hand, OMG-Llava [27] shows better ability in VQA.
The second shortcoming we discuss is their degraded ability to visually ground objects. Surprisingly, although they were trained with pixel-level grounding supervision, not all of these models show superior grounding performance. Figure 3 shows the second prompt to generate a segmentation mask for the ground-truth referring expression. The purpose of this probing is to understand whether the failure in these models is purely on the VQA task, or its inability to ground the objects of interest in the corresponding question or both. Figure 3 shows the worst two models in this aspect, which are GLAMM, the region captioning variant, and Llava-G. Both fail to segment the specific object in question, while OMG-Llava shows better performance.
@@
Third, we highlight another shortcoming, where these MLLMs exhibit degraded ability to follow instructions. In order to probe this, we use the following prompt: “IMGQUESTION? a.OPTION1 b.OPTION2… Answer with the option’s letter from the given.” Figure 4 shows an example with the answers from the worst two models in this aspect which are LISA [9] and Llava-G [26]. Both are incapable of following the instruction, yet Llava-G tries to tackle the question unlike LISA. On the other hand, OMG-Llava shows better ability to follow the instruction and answer the question.
@@
Baselines and upper bounds. In addition to evaluating state-of-the-art pixel-level MLLMs, we propose two baselines and one upper bound. The first of which is inspired by a concurrent work [2] that identified the emergent grounding in multi-modal large language models without the need for any pixel-level grounding supervision. Specifically, we use their attend and segment meta architecture as one of our baselines. However, we are the first to discuss when does such grounding emerge in these models. We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights on how these models reason.
@@
The attend and segment meta-architecture extracts the raw attention map for the output token, , where are the number of layers and heads, resp. Then, are the number of input language tokens before and after the visual tokens respectively, while are the height and width of the input image. Only the attention corresponding to the visual tokens of length are used, and these attention maps are averaged across the layers and heads, resulting in . This is further normalized across all the output, for output tokens. The attend and segment depends on using spaCy natural language processing tool [6] to identify the noun phrases and associate them to the ground-truth referring expressions. Thus, the spaCy embeddings closest to the ground-truth expression are used in the mask selection. This is followed by extracting the maximum attention point to feed into SAM [8] as a point prompt.
For our baseline and upper bound, we build upon the previous pipeline and build an oracle upper bound and an automatic baseline. We introduce two main modifications to account for our observation that the correct grounding can occur with different output tokens describing the object not necessarily aligning with the exact ground-truth expression. The first modification is to go through all the potential output tokens without relying on spaCy embeddings. In the oracle we rely on the ground-truth mask to select the correct token and its corresponding attention map with highest intersection over union as an upper bound. The automatic baseline uses a simple but powerful mechanism where we overlay the predicted masks on the original image to highlight the potential object of interest. This is followed by feeding these images to a multi-modal LLM inquiring on which is best in highlighting this object. Specifically, we use the following prompt “Select the image that has EXPR best highlighted in red color than the others? Answer with a number from 1 to N and mention the number only. IMG”, where EXPR and IMG are the ground-truth expression and the image tokens respectively. In our automatic baseline we rely on GPT-4o for the mask selection. The second modification, since SAM has a good understanding of point prompting ambiguity, we process three potential output masks for each prompt instead of one only. This enables us to utilize the power of SAM in identifying fine-grained objects and referring expressions that tends to surpass what other MLLMs do, even those trained with pixel-level grounding supervision.
| Method | PixGr. | PixMMVP | PixCV-Bench | ||||||||
| Llava 1.5 (7B) [11] | ✗ | 27.3 | 28.0 | - | - | - | 17.4 | 60.3 | - | - | - |
| Llava 1.5 (13B) [11] | ✗ | 39.3 | 30 | - | - | - | 14.5 | 61.4 | - | - | - |
| Cambrian (8B)* [18] | ✗ | 52.0 | 52.0 | - | - | - | 62.2 | 72.2 | - | - | - |
| OMG Llava (7B)** [27] | ✓ | 12.0 | 12.0 | 17.8 | 38.0 | 18.2 | 12.0 | 42.1 | - | 50.5 | 45.9 |
| GLAMM (7B) [16] | ✓ | 1.3 | 2.7 | 31.5 | 47.4 | 5.1 | - | - | 30.2 | 51.9 | - |
| GLAMM - RegCap (7B) [16] | ✓ | 12.7 | 6.7 | 14.5 | 18.6 | 15.1 | 27.8 | 54.4 | 3.6 | 7.4 | 13.0 |
| LISA (7B) [9] | ✓ | 7.3 | - | 18.1 | 42.9 | 12.5 | 3.7 | - | 16.8 | 48.1 | 6.7 |
| Llava-G (7B) [26] | ✓ | 9.3 | - | 17.8 | 13.5 | 12.2 | 14.1 | 4.4 | 1.7 | 17.6 | 15.8 |
| Llava 1.5 (7B) + (a+s) [2] | ✗ | 27.3 | 28.0 | 11.1 | 11.2 | 16.0 | 17.4 | 60.3 | 5.2 | 15.7 | 24.9 |
| Llava 1.5 (13B) + (a+s) [2] | ✗ | 39.3 | 30 | 9.8 | 11.4 | 17.7 | 14.5 | 61.4 | 4.7 | 14.9 | 24.0 |
| Cambrian (8B)* + (a+s) [2] | ✗ | 52.0 | 52.0 | 14.3 | 15.1 | 23.4 | 62.2 | 72.2 | 18.6 | 15.9 | 29.6 |
| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 18.8 | 25.9 | 26.9 | 17.4 | 60.3 | 5.4 | 28.5 | 38.7 |
| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 16.9 | 25.0 | 30.6 | 14.5 | 61.4 | 4.8 | 27.6 | 38.1 |
| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 29.6 | 30.3 | 38.3 | 62.2 | 72.2 | 23.9 | 33.1 | 45.4 |
| Upper Bound - Oracle Selection | |||||||||||
| PixFoundation (7B) (Ours) | ✗ | 27.3 | 28.0 | 26.1 | 38.0 | 32.2 | 17.4 | 60.3 | 6.3 | 49.7 | 54.5 |
| PixFoundation (13B) (Ours) | ✗ | 39.3 | 30 | 23.6 | 38.2 | 38.7 | 14.5 | 61.4 | 5.3 | 50.6 | 55.5 |
| PixFoundation (8B)* (Ours) | ✗ | 52.0 | 52.0 | 52.0 | 56.1 | 54.0 | 62.2 | 72.2 | 54.3 | 64.4 | 68.1 |

### v4

@@ Abstract
Multiple works have emerged to push the boundaries of multi-modal large language models (MLLMs) towards pixel-level understanding. The current trend is to train MLLMs with pixel-level grounding supervision in terms of masks on large-scale labelled data and specialized decoders for the segmentation task. However, we show that such MLLMs when evaluated on recent challenging vision-centric benchmarks, exhibit a weak ability in visual question answering (VQA). Surprisingly, some of these methods even downgrade the grounding ability of MLLMs that were never trained with such pixel-level supervision. In this work, we propose two novel challenging benchmarks with paired evaluation for both VQA and grounding. We demonstrate that simple baselines that are not unified achieve performance that matches or surpasses some of the pixel-level MLLMs. Our paired benchmarks and evaluation enable additional analysis on the reasons for failure with respect to VQA and/or grounding. Furthermore, we propose a prompt sensitivity analysis on both the language and visual prompts tailored for the grounding task. More importantly, we study the research question of “When does grounding emerge in MLLMs with respect to the output tokens?” We propose an interpretability tool that can be plugged into any MLLM to study the aforementioned question. We show that grounding does not necessarily coincide with the exact referring expression in the output, but can coincide with the object parts, its location, appearance, context or state. Code and datasets are publicly available at https://msiam.github.io/PixFoundationSeries/.
@@ 1 Introduction
There have been numerous advancements in pixel-level understanding, including semantic, instance, panoptic segmentation and their video counterpart [36, 18, 11, 21]. In addition, significant progress has been made in visual grounding and reasoning [20, 12], depth estimation [30] and pixel-level tracking [27]. Visual grounding is tied to relating the visual content with language; examples of its pixel-level tasks include referring segmentation and grounded conversation generation. The majority of these models and tasks have been transformed with the emergence of foundation models [2], especially multi-modal large language models (MLLMs) [15, 5]. Nonetheless, pixel-level MLLMs, i.e., MLLMs designed for pixel-level tasks, have shown degradation in their language capabilities [12].
@@
Recent efforts explored the shortcomings of standard MLLMs in vision-centric benchmarks [24, 23]. Such benchmarks focused on challenging visual tasks that require a form of grounding, such as counting. Nonetheless, these benchmarks did not evaluate the recent pixel-level ones. In this work, we focus on pixel-level visual grounding with a segmentation mask output and propose challenging vision-centric benchmarks that are dedicated to evaluating pixel-level MLLMs, which we call PixMMVP and PixCV-Bench. We then provide a comprehensive paired evaluation for VQA and grounding. Our paired evaluation means that the referring segmentation is related to the object of interest in the visual question. Through these, we answer the first research question; “Are the current pixel-level MLLMs trained with mask supervision heading in the right direction to improve both grounding and visual question answering (VQA)?”. Our findings show that the majority of pixel-level MLLMs still fall short in such a challenging setting. While evidently, some of these show superior performance in grounding, we show that such MLLMs with specialized segmentation decoders in their architecture and with mask supervision can be worse than their counterpart baselines when considering both visual and language abilities.
There has been concurrent work [29] that observed the degradation of pixel-level MLLMs’ VQA abilities. Nonetheless, previous efforts used standard evaluation benchmarks that evaluate each task separately. Our benchmarks provide a paired evaluation designed to be vision-centric, with a focus on what MLLMs fall short in and provide the means to interpret the failures of these MLLMs and whether they are stemming from grounding, language capabilities or both. More importantly, unlike concurrent efforts, we focus on the second question of “When does grounding emerge in MLLMs with respect to the output tokens?” Our work documents that emerging grounding in MLLMs does not necessarily coincide with the exact language tokens of the referred object, as shown in Fig. 1. We design an interpretability tool that uses the maximum attention point with respect to the output tokens and analyze the ones that correspond to the best emergent point grounding, using quantitative and qualitative means. While recent works show training-free segmentation emerging from vision language models corresponding to the referred expression [25, 16, 7, 3], our study quantifies that, within MLLMs, that is not necessarily the case.
@@
In summary, our contributions include: (i) Proposing paired pixel-level vision-centric benchmarks, PixMMVP and PixCV-Bench, with segmentation annotations and referring expression of the object of interest in the corresponding questions. In addition to proposing prompt sensitivity variations tailored to the grounding task. (ii) Benchmarking recent efforts in pixel-level MLLMs where we show that the majority degrade VQA capabilities. More importantly, some of them lag in visual grounding with respect to baselines that are not unified. (iii) We propose a novel interpretability mechanism to study when grounding emerges in MLLMs with respect to the output tokens that benefit from our paired benchmarks. Our mechanism uses the observation that grounding can emerge corresponding to different output tokens describing the object’s appearance or location, not necessarily the exact text of the object of interest.
@@ 2 Related work
Pixel-level vision foundation models. There have been various vision foundation models trained for the segmentation task (e.g., SAM and SAM 2.0) [11, 21]. Orthogonal to this, some methods discussed the ability of vision foundation models such as CLIP and BLIP in image segmentation without any segmentation supervision [16, 7, 25]. Yet, they relied on earlier foundation models that did not incorporate the power of large language models. Combining large language models with vision has been extensively researched with pioneering works such as LLaVA [15, 14] and instruct-BLIP [5].
Multiple works afterwards focused on pixel-level visual grounding in these MLLMs with mask supervision and specialized segmentation decoders [12, 20, 33, 34]. However, these methods were lagging in their chat performance. Notably, pixel-level MLLMs were not evaluated on the challenging benchmarks that focused on the shortcomings of MLLMs [24, 23]. Hence, it is still unclear if mask supervision helped to improve their grounding ability on these challenging tasks. In this work, we focus on the previous question to have a better understanding of their performance. While concurrent work [26] designed better pixel-level MLLMs that are capable of strong VQA and grounding, we study it in relation to its baseline standard MLLM that was not trained with mask supervision and compare their robustness within a language and visual prompt sensitivity evaluation.
Benchmarking multi-modal large language models. There is an abundance of standard benchmarks used for evaluating MLLMs (e.g., MMU [32]) and visual grounding benchmarks (e.g., refCOCO/+/g [31, 10]). These have pushed the limits on MLLMs capabilities in terms of VQA and visual grounding. Nonetheless, there have been various works that discussed the shortcomings of MLLMs. One of them discussed the shortcomings in CLIP [19], which is used in various MLLMs as a visual backbone. They proposed a benchmark, MMVP [24], that is focused on the visual aspects within a VQA task. More recently, CV-Bench [23] focused on two major tasks that are vision-focused, which are counting and relative positioning. Both were proposed to evaluate MLLMs that do not have the ability to generate segmentation output. Nonetheless, they provide quite challenging scenarios that can act as a strong benchmark for pixel-level MLLMs.
In this work, we extend these two benchmarks with segmentation annotations and referring expressions that correspond to the object of interest within the VQA task, and propose a paired evaluation metric. Moreover, we evaluate these models within a prompt sensitivity evaluation for both language and visual prompts following previous works [22, 4]. We further introduce novel visual prompt variations specifically designed for the grounding task, pushing the limits of pixel-level MLLMs.
@@ 3 Method and benchmarks
In this section, we describe our two benchmarks and probing techniques for pixel-level MLLMs and MLLMs that were not trained with mask supervision. We focus on the referring segmentation as a representative task of visual grounding. Furthermore, we use visual question answering as a representative task of their vision-language capabilities, with emphasis on the language.
3.1 Paired benchmarks for VQA and grounding
@@
PixMMVP benchmark. We build upon the recently released MMVP [24], which identified clip blind pairs and used them to build a challenging benchmark with the corresponding questions and choices for 300 images. We manually annotate each question with the corresponding object of interest referring expression. There are seven questions only that are not designed to inquire about a specific object in the scene, which are excluded, such as questions inquiring on the view direction of the camera. The referring expressions in our dataset correspond to what fine-grained parts need to be grounded in the image to answer the question. Afterwards, we manually label these objects of interest with polygonal annotations using the VGG annotator [6]. Hence, we create the first paired benchmark for both VQA and pixel-level visual grounding.
@@
PixCV-Bench benchmark. For this benchmark we build upon the 2D component of the recently released CV-Bench [23]. We specifically select the 2D component, since they are sourced from segmentation datasets (i.e., ADE20K [35] and COCO [13]), which can be used in our proposed benchmark. However, the publicly released CV-Bench does not identify the objects in question and their corresponding segmentation. As such we use GPT-4o to parse the questions and identify the objects of interest automatically, followed by manual inspection and correction. This provides us with the categories per question that highlight the objects of interest. While seemingly these are categorical annotations, not referring expressions, certain scenarios in CV-Bench are different. Specifically, in the relative positioning task, all the questions that include an object highlighted by a red box in the image are annotated with the referring expression, “(annotated by the red box)”, beyond simple categories.
@@
Afterwards, we use the expressions from GPT-4o to retrieve the corresponding segmentation mask per image. Furthermore, we use a custom annotation tool to manually filter the objects in question, e.g. selecting only the object mask annotated by the red box and filtering out other instances. Another example that needs manual filtration is when the class in question is a broader category than what is inquired about. Moreover, we identify missing annotations and manually annotate these missing objects. We provide the final paired PixCV-Bench with referring expressions, their segmentation annotations, visual questions and corresponding answers. The supplementary provides visual examples from our benchmarks.
3.2 Pixel-level MLLMs study
We utilize the two proposed benchmarks, PixMMVP and PixCV-Bench, and perform an additional prompt sensitivity analysis to evaluate and inspect the failures of these pixel-level MLLMs. The goal is to push these models in the direction of learning visual grounding abilities on the pixel-level without sacrificing the language abilities that are integral to grounding.
@@ 3.2 Pixel-level MLLMs study
Pixel-level MLLMs failures. We highlight the failures of the current state-of-the-art pixel-level MLLMs through three probing techniques. First, we highlight the degraded performance in VQA from most of these MLLMs that are trained with mask supervision. We use the following prompt, “IMGQUESTION? OPTION1 OPTION2…”, as shown in Fig. 2a. Notably, the worst models in this task, e.g., GLAMM [20], are not able to provide an answer and rather refer to a segmentation mask. On the other hand, OMG-LLaVA [34] tries to tackle the question, even if incorrectly.
@@
The second failure we discuss is that these MLLMs exhibit a degraded ability to follow instructions. In order to probe this, we use the following prompt: “IMGQUESTION? a.OPTION1 b.OPTION2 … Answer with the option’s letter from the given.” Figure 2b shows an example from one of the worst models in this aspect which is LLaVA-G [33] that is incapable of following the instruction, yet tries to tackle the question in free-form. On the other hand, OMG-LLaVA shows better ability to follow the instructions and answer the question.
Third, we highlight their degraded ability to visually ground objects. Surprisingly, although they were trained with mask supervision for grounding, not all of these models show superior grounding performance on our benchmarks. Figure 2c shows the prompt to generate a segmentation mask for the queried referring expression. The purpose of this probing is to understand whether the failure in these models is purely in its language capabilities, or its inability to ground the objects of interest in the corresponding question or both. Figure 2c shows LLaVA-G and OMG-LLaVA, where the latter shows better performance.
Prompt sensitivity evaluation. We propose language and visual prompts variations tailored to our paired benchmark to measure the sensitivity of both pixel-level and standard MLLMs. We argue that language is an integral component in the visual grounding task and, consequently, better language abilities in MLLMs can entail better grounding. Figure 3 shows a Qwen2.5-VL failure from a language prompt variation, where both ask the model to locate the dog’s face, yet the model fails in the second prompt. This motivates our paired evaluation study, which is augmented with a prompt sensitivity evaluation. We conduct the prompt sensitivity evaluation on PixMMVP, since the dataset is considered out-of-distribution with respect to what the majority of MLLMs were instruction-tuned on, unlike PixCV-Bench.
For the language variations, we follow previous works [4] that rely on three variants. These include: (i) spelling errors from omission, transposition, and addition of letters that are randomly generated across eight random locations in the text of the question, the choices and instructions, (ii) random selection of the template for the question and choices and (iii) paraphrasing of the question using GPT-4o. For the visual prompts variations, we propose a novel method that is tailored for visual grounding, where we propose guided random cropping and masking. In the former, the ground-truth mask is used to crop more than 50% of the referred object, while the latter uses a black rectangle to randomly mask out more than 50% of it. Additional details are included in the supplementary.
| Image | Referring Expression | Concept Category | Noun Phrase | Output |
| 1 | the butterfly’s wings | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
| 3 | the flame of the match | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
| 6 | the dog’s face | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
| 161 | the minute hand of the clock | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
1
3
6
161
3.3 Interpretability mechanism
@@
In addition to evaluating state-of-the-art pixel-level MLLMs, we propose a novel interpretability mechanism to study when grounding emerges in MLLMs with respect to the output tokens. We are inspired by a concurrent work [3] that identified the emergent pixel-level grounding in MLLMs without the need for any mask supervision. Specifically, we use their attend and segment meta architecture as one of our baselines. However, we are the first to discuss when such grounding emerges in these models. We identify an interesting connection between the identified output tokens and the output grounding from the attention maps that gives insights into how these models reason.
@@
We extract the raw attention map for the output token, , where are the number of layers and heads, resp. Then, are the number of input language tokens before and after the visual tokens, respectively, while are the height and width of the input image features. Only the attention corresponding to the visual tokens of length is used, and these attention maps are averaged across the layers and heads, resulting in . This is further normalized across all the output, for output tokens. The attend and segment baseline depends on using the spaCy natural language processing tool [8] to identify the noun phrases and associate them with the queried referring expressions. Thus, the spaCy embeddings closest to the queried expression are used in the token selection and their respective attention maps, , are averaged. This is followed by extracting the maximum attention point to feed into SAM [11] as a point prompt.